Determine the friction force for following cases
a) P = 100 N
b) P = 0 N

And the force P required to start the motion up the incline from rest
1
Expert's answer
2020-08-12T16:17:51-0400

The friction force


"F_f=\mu N=\mu mg\cos\alpha"


Assume that "\mu=0.3" and "P" is the weight of the body. We get


a) "F_f=\mu N=0.3\cdot100=30(N)"


b) "F_f=\mu N=0.3\cdot0=0"


The force P required to start the motion up the incline from rest


"P>mg\sin\alpha+F_f"